OpenStudy (anonymous):

Can some one help me with a fairly lengthy integration question? Compute the intergal of 1/(5-4x-x^2)^(5/2)...I have completed the square and been able to factor out the 9^(5/2) to leave the denominator as ((x-2)^2-1)^(5/2). My solution set tells me to substitute (x-2) as a sin function and I'm totally lost from there

OpenStudy (anonymous):

\[x-2=\sec u~~\Rightarrow~~dx=\sec u\tan u~du\] So the integral becomes \[\int\frac{\sec u\tan u~du}{(\sec^2u-1)^{5/2}}\]

hartnn (hartnn):

(5-4x-x^2)^(5/2) is not ((x-2)^2-1)^(5/2)

OpenStudy (anonymous):

^^ I was operating under the assumption it was - shame on me for not checking OP's work :P

hartnn (hartnn):

lol its ok, you would have figured out :P

OpenStudy (anonymous):

\[\begin{align*}5-4x-x^2&=5-(x^2+4x)\\ &=5-(x^2+4x+4-4)\\ &=5-((x+2)^2-4)\\ &=5-(x+2)^2+4\\ &=9-(x+2)^2 \end{align*}\] You can't just factor out the 9 like that...

hartnn (hartnn):

are you sure your set tells x-2 as sin ? or x+2 as sin ?

OpenStudy (anonymous):

\[\int\frac{dx}{\left(9-(x+2)^2\right)^{5/2}}\] The substitution would then be \[x+2=3\sin u~~\Rightarrow~~dx=3\cos u~du\] and so \[\int\frac{3\cos u~du}{\left(9-(3\sin u)^2\right)^{5/2}}\] NOW you can factor out the 9: \[\int\frac{3\cos u~du}{\left(9-9\sin^2 u\right)^{5/2}}\\ \int\frac{3\cos u~du}{9^{5/2}\left(1-\sin^2 u\right)^{5/2}}\\ \]

OpenStudy (anonymous):

I can't see the sign change when I completing the square...

hartnn (hartnn):

show your work, we'll help you find the error

OpenStudy (anonymous):

I just saw it. I was convinced that was wrong on the solution set. At that point it tells me to set x+2=3sin theta. Where is the 3 coming from?

hartnn (hartnn):

for \(\Large a^2-x^2\) we plug in x = a sin theta

OpenStudy (anonymous):

I can see it now. Thank you. I've been staring at this thing for a couple days and could not find the mistake!

hartnn (hartnn):

glad we could help ^_^ welcome :)

OpenStudy (anonymous):

I've got another thats bugging me if you are still here....

OpenStudy (anonymous):

\[\int\limits_{1}^{4}\sqrt{t}\ln t dt\]

hartnn (hartnn):

i am here :)

hartnn (hartnn):

integration by part

OpenStudy (anonymous):

Let me show you where I'm stuck\[= \ln t \times2/3t ^{(3/2})-\int\limits_{1}^{4}2/3t ^{(3/2)}*(1/t)*t ^{(1/2)}dt\]

OpenStudy (anonymous):

My solution set shows that it should be \[\int\limits_{1}^{4} (2/3) t ^{(3/2)}*t ^{-1}dt\]

hartnn (hartnn):

how did u get that extra t^(1/2) ?

hartnn (hartnn):

\(= \ln t \times2/3t ^{(3/2)}-\int\limits_{1}^{4}(2/3)t ^{(3/2)}\times (1/t)dt\) and 1/t is \(t^{-1}\)

OpenStudy (anonymous):

I am wondering the same thing...

OpenStudy (anonymous):

The formula for integration by parts says \[\int\limits_{a}^{b}uv \prime=uv -\int\limits_{a}^{b}u primev\]

OpenStudy (anonymous):

Ive got u=lnt and u prime =1/t v prime sqrt v and v=2/3(t)^3/2

hartnn (hartnn):

correct

OpenStudy (anonymous):

Oh dear... I just my mistake....its the end of the quarter and I'm fried!!!! Lol. Just typing it and looking at it with you helped me. THANKS!!!

OpenStudy (anonymous):

I just randomly added that t^(1/2). WOW!!!!

hartnn (hartnn):

lol yeah, ans welcome :)